Wesen Corp. will pay a dividend of $4.10 next year. The company has stated that it will maintain a constant growth rate of 5.25 percent a year forever If you want a return of 18 percent, how much will you pay for the stock? (Do not round intermediate calculations and round your answer to 2 declmal places, e.g., 32 Current stock price If you want a return of 10 percent, how much will you pay for the stock? Do not round intermediate calculations and round your answer to 2 decimal places . g., 321 Current stock proe s
Answer #1
Calculation of price of stock when return is 18%
D1= $4.10
growth= 5.25%
Required return= 18%
Value of stock= D1/(required return-growth)
                            =4.10/(0.18-0.0525)
                           =4.10/0.1275
                           =$32.16
Value of stock is $32.16
Calculation of price of stock when return is 10%
D1= $4.10
growth= 5.25%
Required return= 10%
Value of stock= D1/(required return-growth)
                            =4.10/(0.10-0.0525)
                           =4.10/0.0475
                           =86.32
Value of stock is $86.32